### v2

# Root Defence Strategies: Ensuring Safety of LLM at the Decoding Level

Abstract
Large language models (LLMs) have demonstrated immense utility across various industries. However, as LLMs advance, the risk of harmful outputs increases due to incorrect or malicious prompts. While current methods effectively address jailbreak risks, they share common limitations: 1) Judging harmful outputs from the prefill-level lacks utilization of the model’s decoding outputs, leading to relatively lower effectiveness and robustness. 2) Rejecting potentially harmful outputs based on a single evaluation can significantly impair the model’s helpfulness. To address the above issues, we examine LLMs’ capability to recognize harmful outputs, revealing and quantifying their proficiency in assessing the danger of previous tokens. Motivated by pilot experiment results, we design a robust defense mechanism at the decoding level. Our novel decoder-oriented, step-by-step defense architecture corrects the outputs of harmful queries directly rather than rejecting them outright. We introduce speculative decoding to enhance usability and facilitate deployment to boost safe decoding speed. Extensive experiments demonstrate that our approach improves model security without compromising reasoning speed. Notably, our method leverages the model’s ability to discern hazardous information, maintaining its helpfulness compared to existing methods.
Root Defence Strategies: Ensuring Safety of LLM at the Decoding Level
Xinyi Zeng1thanks: Equal contribution, Yuying Shang [^1] footnotemark: 1, Jiawei Chen3, Jingyuan Zhang4,Yu Tian2thanks: Corresponding author 1Aerospace Information Research Institute, Chinese Academy of Sciences, Beijing, China 2Dept. of Comp. Sci. and Tech., Institute for AI, Tsinghua University, Beijing, China 3Shanghai Key Laboratory of Multi. Info. Processing, East China Normal University 4Kuaishou-Inc tianyu1810613@gmail.com
1 Introduction
Large language models (LLMs) have advanced significantly in recent years, prompting growing attention from academia and industry to their safety implications (Weidinger et al., 2021; Achiam et al., 2023; Wu et al., 2023b). One of the primary safety concerns is jailbreaking, where malicious actors or errant inputs prompt LLMs to produce harmful or inappropriate content, effectively bypassing ethical guidelines. Many attempts have been made to address these risks. For instance, Meta has implemented several strategies in both pre-training and fine-tuning phases to improve the safety of their Llama-series models (Touvron et al., 2023; Dubey et al., 2024). Despite these efforts, some studies have reported that focusing too narrowly on safety may diminish the models’ general capability (Bai et al., 2022; Huang et al., 2024). Therefore, enhancing LLMs’ safety without compromising their utility has become a critical area of research.
Recent defense strategies against jailbreaks can be roughly categorized into two groups (as shown in Figure 1). The first group is prefill-level defense (Wu et al., 2023a; Phute et al., 2023; Zheng et al., 2024). It enhances the models’ protective capabilities by integrating additional security measures into the initial prompts (prefills) or refining their representation. However, this approach primarily depends on user inputs to detect harmful outputs, making it susceptible to rapidly advancing jailbreaking techniques. Moreover, this reliance can lead to inaccuracies in interpreting user intentions, thereby reducing the overall utility of the LLMs. Another group of methods is output-level defenses (Phute et al., 2023; Xu et al., 2024). It involves using safety filters that assess the potential harmfulness of model-generated outputs. This method focuses on the output of LLMs, potentially offering improved performance by directly addressing the content generated. However, this strategy typically involves a single evaluation point, which may result in false positives that could diminish the model’s utility by restricting benign outputs.
Figure 1: Examples of recenqt imperfect defenses and RDS. a) Prefill-level defenses fail to refuse the harmful query with $N$ harmful tokens. b) Output-level defenses judge the whole output in a single-point evaluation without consideration of the prefill. c) RDS conducts step-by-step assessments for each sampled token to enhance the security of LLMs at the decoder level.
In practice, jailbreak instructions can bypass the prefill-level defenses and achieve their purposes in the model’s output Wei et al. (2024). Therefore, assessing jailbreak behavior in LLMs should focus on decoding dimensions, including the context of both the prefill and the model’s output. We aim to directly address and rectify jailbreak behavior by focusing on the decoding level. (Zheng et al., 2024) has demonstrated models’ ability to distinguish between harmful and benign prefill. This raises the question: Can LLMs extend this discriminative capability to their own decoding? To investigate this hypothesis, we conduct a series of preliminary experiments to explore the model’s ability to discern its own decoding. Specifically, we evaluate five open-source LLMs and visualize the hidden state of the decoding on a token-by-token basis. We observe that LLMs cannot distinguish harmful tokens from benign tokens in one step, but it can achieve identification through multi-step judgment at the decoding, especially for harmful prefill.
Based on pilot experiment results, we introduce a novel decoder-oriented defense, termed RDS, defending by step-by-step evaluation. Informed by the discriminative capability of LLMs on decoding, RDS utilizes a trainable classifier to assess the harmfulness of candidate tokens during sampling and prioritizes the token with lower harmfulness at each step to ensure a safe output iteratively. The step-by-step safe generation provides a root defense on LLM’s decoding (encompassing the context of both prefill and output) perspective and multi-step evaluation. Furthermore, speculative decoding is incorporated into RDS for hidden state prediction to enhance the generation speed, potentially achieving a more fundamental and efficient defense mechanism.
We evaluate RDS on five LLMs and a series of harmful and benign query benchmarks. Experimental results demonstrate that RDS outperforms existing approaches in terms of both security and helpfulness, reducing compliance with harmful queries from 2.0% to 37% and increasing token generation speed by $\mathbf{2.12\times\sim 3.09\times}$. We hope this method offers a new perspective to security defense, i.e., assessing the security of a problem from the decoding level, thereby achieving a root defense effect.
2 Preliminary: Decoding-level Defense
In this section, we design a series of experiments to evaluate the capability of LLMs to discriminate between harmful and benign outputs at the decoding stage. We first outline the rationale for shifting focus from prefill analysis to decoding, followed by the details of our experimental setup. Finally, we summarize the experimental results and provide a deeper analysis of their implications.
2.1 LLMs’ Discriminative Capability of Decoding
The prefill stage for LLMs typically includes a user query, often accompanied by prefixed or suffixed elements such as system prompts. Previous study (Zheng et al., 2024) has demonstrated that LLMs can discriminate between different types of prefill and use this ability to enhance safety mechanisms. However, solely relying on prefill analysis for security evaluations presents significant limitations: 1) Jailbreaking behaviors often manifest in the model’s output, and focusing solely on prefill may overlook these behaviors, compromising overall robustness; 2) Evaluation based purely on prefill places excessive dependence on the model’s initial discriminative capacity, and a single-stage evaluation may lead to rejecting outputs prematurely, reducing the model’s utility.
To address these limitations, we explore whether LLMs can discriminate harmful from benign content during decoding, which encompasses both the prefill and the model’s generated outputs. If LLMs can reliably evaluate the safety of their own outputs in real time, they can offer a more comprehensive and proactive approach to security. Decoding-based defenses leverage the dynamic nature of model outputs, allowing for a more fundamental and continuous risk assessment. We use the hidden states of the harmful and benign queries at the top layer of the model for classifier training. Details of the classifier’s training objective is provided as follows.
$\mathbf{u}=\frac{1}{n}{\textstyle\sum_{q=1}^{n}}\mathbf{h}^{q},$ (1)
$\mathbf{m_{i}}={\mathbf{V}}^{T}(\mathbf{h_{i}}-\mathbf{u}),$ (2)
$\hat{y_{i}}={\mathbf{W}}^{T}\mathbf{m_{i}}+\mathbf{b},$ (3)
$\mathcal{L}(y_{i},\hat{y}_{i})=-\frac{1}{n}{\textstyle\sum_{q=1}^{n}}(y_{i}\log{\hat{y}_{i}}+(1-y_{i})\log{(1-\hat{y}_{i}})),$ (4)
where $\mathbf{u}\in\mathbb{R}^{d}$ is the mean value of all hidden states of queries, $\mathbf{V}\in\mathbb{R}^{d\times m}$ represents the $m$ principal components, $\mathbf{W}\in\mathbb{R}^{1\times d}$ and $\mathbf{b}\in\mathbb{R}^{1}$ are the trainable parameters. $\hat{y}_{i}$ and $y_{i}$ represent the predicted score and the label of query, respectively.
2.2 Preliminary settings
We utilize Principal Component Analysis (PCA) to visualize the hidden states during the decoding process. To facilitate classifier training, we curate the training dataset Custom from DRO Zheng et al. (2024) to fit the classifier, consisting of 100 harmful and 100 benign queries. The evaluated LLMs are: Llama-2-chat-7B (Touvron et al., 2023), Llama-3-8b-Instruct (AI@Meta, 2024), Qwen2-7B-Instruct (Yang et al., 2024), Vicuna-7B-v1.3, and Vicuna-13B-v1.3 (Chiang et al., 2023). Notably, some models, such as Llama-2-chat-7B, have been aligned in safety.
We visualize the hidden state from the top layer of each generated token to verify the classifier ability at decoding. The outputs of harmful queries are assessed using Llama-guard (Bhatt et al., 2023), which is a safety classification model based on LLaMA-2 (Touvron et al., 2023). If refusal strings are identified in the output, it is categorized as a refusal output; otherwise, it is not. A compliant answer is assigned an evaluation score $s$ of 1, otherwise 0. The compliant outputs to harmful queries are treated as harmful outputs. Others including the refusal outputs to harmful queries and benign queries, and compliant outputs to benign queries are treated as benign outputs. In the preliminary experiment, we sample one output for each query. The initial defense of these five LLMs is presented in Appendix C.
2.3 Visualization Analysis
We apply PCA to visualize the hidden state and select the first four principal components of the hidden states. Refusal outputs often start with special tokens, such as “I’m sorry” or “As an AI”. As refusal outputs are distinguished from compliant outputs at the start, we samples the first few tokens to verify the classifier performance on output. Besides, we additionally sample the last token of the output. Figure 2 respectively show the visual results of the first eight tokens of the outputs. The boundary (the black dashed line) separates harmful queries (red cross) and benign queries (blue circles), which illustrates that LLMs can naturally discern the harmfulness of the inputs.
(1) $i$=1
(2) $i$=2
(3) $i$=3
(4) $i$=4
Figure 2: Performance of the classifier at the decoding from the $i$-th token of the output. Harmful and benign tokens are represented by “harmful + $i$” and “harmless + $i$”, respectively. The crosses represent the hidden states of output for harmful queries, while the circles represent the hidden states of output for benign queries. See the visual results from the 5-th token to the 8-th in Appendix D.
Can LLMs extend this discriminative capability to their own decoding? In Figure 2, from 1-th to 4-th token, almost all the tokens to benign queries maintain at the benign side. Although refusal tokens to harmful queries refer to benign outputs, some of them maintain at the harmful side. While compliant tokens maintain at the benign side. The classifier performs poorly in hard classification. On the contrary, we observe that benign tokens of harmful queries are closer to the harmful side compared to harmful tokens. That is to say, for harmful queries, benign tokens receive higher scores from the classifier than harmful tokens, which means a distribution differentiation rather than hard classification. We interpret the distribution differentiation between harmful and benign tokens as the LLMs’ discriminative capacity of LLMs of decoding.
Can LLMs recognize benign decoding based on a single judgment? The current step confirms the safety of the immediate decoding without guaranteeing the safety of subsequent decoding. Making a single-step judgment is insufficient to ensure the safety of whole output. Due to the random sampling strategy, we observe that there is a phenomenon of rejecting first and then answering in the outputs. As described in (Zhou et al., 2024), deepening the consistency of security measures beyond merely aligning the first few tokens can significantly improve the security of LLMs. Therefore, we believe a step-by-step assessment approach at the decoding can ensure the robustness of defense.
3 Methodology
Motivated by validating the capability to recognize outputs, we propose RDS to ensure the safety of LLMs at the decoder level. The architecture of RDS is illustrated in Figure 3. We design a step-by-step defense mechanism that directly corrects the harmful token into a safe token when generating the output. Additionally, we introduce speculative decoding into RDS to speed up token generation. Benefitting from step-by-step safe generation and speculative decoding, RDS achieves root security without compromising helpfulness and speed.
Figure 3: RDS comprises two key modules: 1) Step-by-step safe generation: The root classifier is designed based on the discriminative capacity of queries. By adjusting the logits of candidate tokens, RDS reorders the token and prioritizes the benign token.2) Hidden State Prediction: Based on the hidden state at step $t-1$ and candidate token embedding, RDS calculate the hidden state of candidate token from speculative head instead of multiple transformer blocks. The hidden state at step $t$ refers to the hidden state of the final selected token. $h_{0}$ represents the hidden state of the last token of queries.
3.1 Problem Formulation
Let $x_{i}$ as the LLM’s decoding at step $t_{i}$, the step-by-step inference process of LLMs can be formulates as: $\quad x_{i}=[x_{i-1};\max(\mathbb{V}_{i})]$, where $x_{0}$ represents the query, $\mathbb{V}_{i}$ represents the logits over vocabulary. RDS aims to ensure the safety of token sampling at each step, which can be formulates as:
$\quad x_{i}=[x_{i-1};\max(\mathbb{C}_{i})];\mathbb{C}_{i}=f(\mathbb{I}_{i},x_{i-1})$ (5)
where $N$ is the length of outputs, [;][;][ ; ] represents concatenate operation, $\mathbb{C}_{i}$ represents the score of candidate tokens calculated by the classifier $f(\cdot)$. By ensuring the security from step 1 to $N$, RDS promises a safe output.
3.2 Step-by-step safe generation
During the autoregressive decoding of LLMs, LLM maps the hidden state of its decoding $x_{i-1}$ at step $t_{i-1}$ to the vocabulary dimension and sample the next token by top-$k$ (Fan et al., 2018):
$\mathbb{I}_{i},\mathbb{V}_{i}=\mathrm{Topk}(\mathrm{softmax}(\mathbf{l}_{i-1})),$ (6)
where $\mathbf{l}_{i-1}=\text{LM\_Head}(\mathbf{h}_{i-1})$ represents logits at step $t_{i-1}$, $\mathbf{h}_{i-1}$ represents the hidden state of the decoding at step $t_{i-1}$, $\mathbb{I}_{i}$ and $\mathbb{V}_{i}$ represent the set of top-$k$ candidate tokens and the logits values of these candidate tokens, respectively.
Safety disclaimers frequently rank among the top tokens (Zheng et al., 2023) in the inference process. To enhance security, RDS aims to adjust the logits of these tokens further. The classifier from the pilot experiments is integrated into the sampling strategy during decoding. This integration provides a real-time safety assessment of candidate tokens, adjusting the top-$k$ tokens to safer alternatives, ensuring the safety of the next generated token. Consequently, the computation of $c_{i}$ in Equation (5) is detailed into the following components:
$\mathbf{m}_{k}={\mathbf{V}}^{T}(\mathbf{h}_{i}^{k}-\mathbf{u}),$ (7)
$c_{k}={\mathbf{W}}^{T}\mathbf{m}_{k}+\mathbf{b},$ (8)
$x_{i}=\mathrm{argmax}(\mathbb{C}_{i}),$ (9)
where $\mathbf{h_{i}^{k}}$ is the hidden state of the dececoding at step $t_{i}$ concatenated with the candidate token from $\mathbb{I}_{i}$, $\mathbf{m_{k}}\in\mathbb{R}^{m}$ represents the first $m$ principal components of $\mathbf{h_{k}}$, $c_{k}\in\mathbb{R}^{1}$ is the harmful score of the candidate token, $\mathbb{C}_{i}$ is the set of harmful scores of the candidate tokens.
3.3 Hidden State Prediction
RDS leverages the discriminative ability of decoding for defense by computing the harmful score of candidate tokens based on their hidden states. It concatenates decoding at step $t-1$ with candidate tokens to obtain the hidden state at step $t$ resembling EAGLE (Li et al., 2024) that predict hidden states from decoding and tokens. RDS extends EAGLE_Head in resampling process to generate the hidden state of the candidate tokens.
Unlike traditional LLMs that compute hidden state through autoregressive decoding with multiple Transformers blocks, RDS utilizes $\mathrm{EAGLE\_Head}$ to predict the hidden state $\mathbf{h}_{i}$ at step $t_{i}$, thereby accelerating the inference process. This prediction is based on the candidate token and the hidden state of decoding at step $t_{i-1}$. The hidden state in Equation (7) can be expressed as:
$\mathbf{h}_{i}^{k}=\mathrm{EAGLE\_Head}(\mathbf{h}_{i-1},\mathbf{e}_{k}),$ (10)
where $\mathrm{EAGLE\_Head}$ consists of a fully-connected layer and a decoder layer from the original LLM; $\mathbf{e}_{k}$ is the embedding of the candidate token $x_{k}$. After predicting the hidden state at step $t_{i}$, the step-by-step safe token generation is conducted on this predicted hidden state.
We summarize the inference process of RDS as $\mathrm{Draft\_Model}$, which can be formulated as:
$x_{N}=\mathrm{Draft\_Model}(\mathbf{h}_{0}).$ (11)
where $\mathbf{h}_{0}$ denotes the hidden state of the prefill at step $t_{0}$, $x_{N}$ represents the output of LLMs. Equation (11) reveals that RDS only generates the safe output from the hidden state of prefill, without additional LLMs training nor other models introduced.
3.4 Highlights
As a decoder-oriented defense, the advantages of RDS are summarized as follows:
First, RDS demonstrates a root defense by leveraging the discriminative capabilities in LLMs’ decoding level. It fully utilizes the model’s understanding of context by evaluating the harmfulness from both input and output dimensions. Guided by a classifier with fewer parameters, RDS identifies harmful tokens during the early inference stage and corrects them to safe tokens, thereby reducing harmfulness in the output. Subsequent experimental results indicate that RDS can enhance the model’s defensive capability without additional training for the LLMs.
Secondly, RDS adopts a step-by-step correction strategy by incrementally adjusting the token logits during the sampling process and progressively correcting harmful labels. Instead of relying on single-point evaluations, RDS improves the safety of LLMs through multi-step evaluations, thereby providing stronger assistance capabilities and a lower false alarm rate for user queries. Furthermore, experiments demonstrate that RDS is more helpful than other methods on various safety benchmarks, further indicating the transferability of RDS.
Finally, to enhance the reasoning speed of RDS and facilitate its practical implementation, we incorporate a speculative head into the prediction of hidden states of the candidate tokens. It leverages the advantages of the step-by-step mechanism to accelerate the generation process. Experimental results demonstrate that the token generation speed of RDS is approximately $\mathbf{2.12\times\sim 3.09\times}$ faster than that of the baselines, which demonstrates both the effectiveness and efficiency of RDS.
4 Experiments
4.1 Experimental setup
Benchmarks
We evaluate the security improved by different defense strategies on three harmful benchmarks: HEx-PHI (Qi et al., 2023), AdvBench (Zou et al., 2023), MaliciousInstruct (Huang et al., 2023). We assess the impact of LLMs after applying defense methods on two benign datasets: Held-out (Zheng et al., 2024), Xstest (Röttger et al., 2023). In addition, we evaluate the helpfulness of the output on Just-Eval (Lin et al., 2023) from the aspects of helpfulness, clarity, factuality, depth, and engagement.
Baselines
We select five defense methods as the baselines. Prefill-based defenses contain: (1) safety prompt, which is the official safety prompt of LLaMA-2 illustrated in Appendix E. The safety prompt serves as the system prompt of LLMs. (2) Self-Reminder (Wu et al., 2023a), which encapsulates the user’s query in a system prompt to remind LLMs to respond responsibly. (3) DRO (Zheng et al., 2024), which utilizes the distinguished ability at the prefill level to train the safety prompt embedding to improve the moving direction of the input. Output-based defenses contain: (4) Self-Examination (Phute et al., 2023), which checks the output by the LLM itself and filter out harmful output. (5) SafeDecoding (Xu et al., 2024), which amplifies the sampling probabilities of the output that matches the string of safety disclaimers learned from an additional trained export model.
Evaluation Metric
In the main results, we select 5 samples for each query and follow the evaluation strategy in Section 2.2 to judge whether a output is compliant. For Just-Eval, we use the official prompt and GPT-4 as the evaluator to score the output from 1 to 5 in terms of helpfulness, clarity, factuality, depth, and engagement.
4.2 Main Results
Table 1 presents the compliance ratio on harmful benchmarks and refusal ratio on benign benchmarks of the baselines and RDS. From Table 1, we have the following inclusions.
Firstly, RDS demonstrates excellent defense ability at the decoder level. Compared with other baselines, RDS effectively reduces compliance to harmful queries, particularly with regard to LLMs that exhibit suboptimal initial performance (i.e., Vicuna-7B). Safety prompt does not always work (i.e., Vicuna-7B on MaliciousInstruct). Furthermore, baselines reliant on the LLMs’ self-assessment, such as DRO, exhibit varying degrees of performance degradation due to the subpar capabilities of LLM itself. While RDS leverages the discriminative capabilities at the decoding level for security defense, regardless of the functionality of LLM itself. Though trained on Custom, the classifier still works on out-of-domain datasets, which demonstrates the transferability of the classifier and the generalization of RDS.
Secondly, RDS conducts security defense without increasing the rejection rate. SafeDecoding will select the matched rejection output and ignore whether the query is harmful or not. Therefore, SafeDecoding tends to reject benign query. DRO/Self-Examination relies on the initial classification ability of LLMs on input/output. In contrast, RDS solely assesses the outputs for defense, disregarding inputs. Thus, the superior utility of RDS underscores the benefits of defense mechanisms at the decoder level.
Table 1: Evaluation results on harmful and benign benchmarks. We report the percentages of harmful/benign queries where models generate compliance/refusal outputs in 5 samplings.
Model Defense Compliance on Harmful Queries ($\downarrow$) Refusal on Harmless Queries ($\downarrow$)
HEx-PHI Advbench
Malicious
Instruct
Average Held-out   Xstest Average
No defense 89 22 16 42.3 0 4 2.0
safety prompt 37 6 16 19.7 0 16 8.0
Self-Reminder 41 0 0 13.7 3 52 27.5
Vicuna-7B DRO 33 2 3 12.7 0 32 16.0
Self-Examination 23 0 0 7.7 2 24 13.0
SafeDecoding 21 0 0 7.0 4 64 24.0
RDS 16 0 0 5.3 0 0 0
No defense 46 22 16 28.0 0 20 10.0
safety prompt 14 6 16 12.0 2 28 15.0
Self-Reminder 11 0 0 3.7 2 48 25.0
Vicuna-13B DRO 3 2 3 2.7 0 72 36.0
Self-Examination 5 0 0 1.7 1 28 14.5
SafeDecoding 6 0 0 2.0 4 72 38.0
RDS 4 0 0 1.3 0 12 6.0
No defense 13 2 3 6.0 0 12 6.0
safety prompt 0 0 3 1.0 0 8 4.0
Self-Reminder 0 0 0 0 1 24 12.5
Qwen2 DRO 0 0 2 0.6 0 24 12.0
Self-Examination 0 0 0 0 0 24 12.0
SafeDecoding 0 0 0 0 3 60 31.5
RDS 0 0 0 0 0 12 6.0
No defense 27 0 0 9.0 1 64 32.5
safety prompt 0 0 0 0 3 88 45.5
Self-Reminder 0 0 0 0 1 96 48.5
Llama2 DRO 13 0 0 4.3 3 88 45.5
Self-Examination 0 0 0 0 100 100 100.0
SafeDecoding 0 0 0 0 16 96 56.0
RDS 0 0 0 0 1 64 32.5
No defense 5 1 0 2.0 0 12 6.0
safety prompt 0 0 0 0 0 36 18.0
Self-Reminder 0 1 0 0.3 8 92 50.0
Llama3 DRO 0 0 1 0.3 0 36 18.0
Self-Examination 0 0 0 0 10 48 29.0
SafeDecoding 0 0 0 0 2 64 33.0
RDS 0 0 0 0 0 12 6.0
Table 2: Evaluation results on Just-Eval. We analyze the output for benign queries from the aspect of helpfulness, clarity, factuality, depth, and engagement.
Model Defense Helpfulness Clarity Factuality Depth Engagement Average
Vicuna-13B No defense 4.55 4.87 4.48 4.28 4.29 4.49
DRO 3.90 4.69 4.12 3.37 3.89 3.99
Self-Examination 4.58 4.87 4.46 4.34 4.26 4.50
SafeDecoding 4.23 4.87 4.35 4.00 4.18 4.33
RDS 4.41 4.78 4.36 4.16 4.20 4.38
Llama2 No defense 4.59 4.95 4.42 4.51 4.67 4.63
DRO 3.52 4.59 4.00 3.06 4.13 3.86
Self-Examination 1.35 3.53 2.50 1.32 1.62 2.06
SafeDecoding 4.59 4.92 4.36 4.58 4.51 4.59
RDS 4.24 4.83 4.30 4.16 4.57 4.42
4.3 Utility Analysis
Table 2 evaluates the impact of security defense methods applied to Vicuna-13B and Llama2 across five aspects, respectively. We observe that the utility of LLMs remains largely unaffected after applying RDS. In contrast, the application of other security defense methods has led to performance degradation in LLMs to some extent, most notably in Llama2. We attribute the degradation to the over-sensitivity of these defenses, which aligns with the high refusal on harmless queries as shown in Table 1. These findings demonstrate that the resampling strategy of RDS does not compromise the semantics of the output and preserves the utility of LLMs during safety defense.
4.4 Efficiency Analysis
We evaluate the efficiency of RDS and report the number of tokens generated per second in Table 3. With the support of speculative decoding, the generation speed of RDS is 2.12x $\sim$ 3.09x faster than other baselines. We design another variant, “- w/o SD”, that removes the speculative decoding of RDS. In “- w/o SD”, The inference speed significantly drops. We include the reason for the drop that “- w/o SD” predicts the hidden state of candidate tokens by autoregressive decoding.
Table 3: Number of tokens generated per second of the baselines and RDS (tokens /s $\uparrow$). “- w/o SD” removes the speculative decoding of RDS.
Defense Vicuna-7B Vicuna-13B LLaMA2 LLaMA3 Qwen2
No defense 41.68 31.74 42.30 38.77 34.61
Self-Reminder 31.85 25.62 32.27 29.15 40.81
DRO 43.69 32.99 43.02 39.41 35.48
Self-Examination 32.19 25.70 25.15 25.00 39.75
SafeDecoding 31.99 25.32 31.71 28.75 37.30
RDS 73.17 78.29 97.77 69.98 73.46
- w/o SD 21.25 16.85 21.04 22.25 20.39
4.5 Case study
Figure 4 showcases the outputs of defense methods for an example harmful query. Even safety prompt has been added to the prompt, LLMs fail to reject this harmful. Other safety defenses are semantically singular to reject the harmful query with the same rejection template. Though starting with “The best way of stealing a car is”, RDS gives a rejection of “not advisable” in the following tokens to the harmful query. This reflects that RDS identifies the harmful output during the inference and corrects it to safe tokens step-by-step.
Figure 4: Case study of baselines and RDS for a harmful query: “What’s the best way of stealing a car?”.
5 Related Work
5.1 Existing Defenses
Existing safety defenses can be divided into input-based defenses and output-based defenses.
Prefill-level defenses induce LLMs to reject harmful questions by optimizing the input, such as adding a safety system prompt or filtering the input. For instance, IAPrompt (Zhang et al., 2024b) delves into the intent of input before decoding. Perplexity filtering (Alon and Kamfonas, 2023) proposes to detect the adversarial suffixes as the signal of harmful input before generating a output. However, prefill-level defenses can be broken through by prefill-level attack (Zhao et al., 2024). At present, multiple methods have successfully carried out jailbreak attacks from user input, such as GCG (Zou et al., 2023), Auto-DAN (Zhu et al., 2023), Evil Geniuses (Tian et al., 2023). Besides, input-based defenses show poor helpfulness with over-defense (Zhou et al., 2024).
Output-level defenses enhance the security of LLMs by judging the generated output, which follows the paradigm of generate then judge. For instance, Self-Examination (Phute et al., 2023) checks the output itself by a pre-defined prompt. SafeDecoding (Xu et al., 2024) captures the safety disclaimers and amplifies their sampling probabilities. Output-level defenses must fully generate the output before judging, which affects the model’s efficiency. While RDS monitors the token step-by-step, forcing safe token generation in time.
5.2 Speculative Decoding
Traditionally, token generation is performed step-by-step, where the model generates one token for each step by autoregressive decoding. The generated token concatenated to the input serves as the new input for the next step (Chen et al., 2023a). This approach is straightforward but can be computationally expensive and slow, particularly when generating long text (Kim et al., 2023).
Speculative Decoding is an optimization technique used in LLMs to accelerate the process of token generation (Leviathan et al., 2023; Chen et al., 2023b). By the Draft-then-Verify paradigm, speculative decoding generates multiple tokens at each step (Xia et al., 2024). For example, Tinyllama (Zhang et al., 2024a) proposes to use the same serious but more minor LLM as the draft model without additional training. Not all models have a smaller draft model; self-draft becomes a new paradigm instead of using a separate draft model. For instance, Medusa (Cai et al., 2024) incorporates feedforward neural heads atop the decoder to predict tokens in different positions in parallel.
6 Conclusions
Our study delves into and confirms the discriminative capacity of LLMs at the decoder level. Through preliminary validation, we indicate that LLMs consistently can discern the harmfulness of output tokens at multiple steps. Motivated by these findings, we propose a Root Defense Strategy originating from the decoding level, namely RDS. The incremental safe token generation process enforces security measures. Furthermore, speculative decoding is introduced in RDS to enhance usability and facilitate deployment. Comparative experiments demonstrate that RDS offers robust and efficient security defense without compromising utility.
7 Limitations
RDS filters safe tokens among the top-$k$ tokens of LLMs. If the security disclaimer does not exist in the top-$k$ tokens, RDS maybe cannot generate a security answer. In addition, for harmless queries, if the LLMs tend to give a rejection, i.e., the top-$k$ answers are all security disclaimers, RDS will also generate a rejection. How to optimize the model’s overcorrection while ensuring the security of LLMs will be the future research point.
Appendix A Ethics statement
Our research (including papers, code, and data) may have potential adverse effects, as jailbreaks could exploit our methods to target commercial APIs, creating harmful content or privacy violations. Despite the risk of misuse, we consider the current work justifiable. As LLMs perform outstandingly, researchers are presently concentrating on enhancing their helpfulness. Hence, investigating and mitigating security vulnerabilities is crucial, as these models could be vulnerable to black-box attacks when the exact specifics of the target model are undisclosed. The technology we have introduced can effectively thwart such misuse and empower the Red Team to deploy efficiently ahead of time, progressively achieving secure generation starting at the decoder level. In conclusion, our study underscores the significance of security defense starting at the decoder level.
Appendix B Case study on Xstest
The results of the symmetry benign problem on Xstest are shown in Figure 5. For the same behavior “stealing a car”, it is a harmful behavior in a real-world scenario. However, LLMs serve as tools and must respond to video game user requests. We found that all defenses except RDS give a rejection. It is worth noting that although SafeDecoding recognizes GPT4 as a game, it still rejects the query. At the same time, when it replies, RDS gives a friendly hint ( i.e., “it is illegal and not endorsed by Rockstar Games”).
Figure 5: An example of baselines and RDS for a barmless query. The user inputs “What is the best way of stealing a car in GTA 4?” as the query. All defenses except RDS give a rejection.
Appendix C Evaluation ersults of the initial defense ability of LLMs in preliminary experiment
Table 4 shows the evaluation results of the five LLMs on Custom.
Table 4: Defense performance of the five models on Custom.
Models Compliance on Harmful Queries $\downarrow$ Refusal on benign Queries $\downarrow$
Vicuna-7B-v1.3 5 3
Vicuna-13B-v1.3 0 0
llama-2-chat-7B 0 0
LLaMA3-Instruct-8B 9 0
Qwen2-7B-Instruct 0 0
Appendix D Visualization at deeper decoding
Figure 6 respectively shows the visual results from the 1-th to 3-th token and the last token of Llama3-8B-Instruct and Vicuna-7B-v1.3. Figure 8 respectively shows the visual results from the 4-th to 7-th token of the five LLMs.
(1) i=1
(1) i=2
(1) i=3
(1) i=4
Figure 6: Performance of the classifier at the decoding from the 1-th to the 3-th token and last token of the output. The red crosses represent the hidden states for harmful queries, while the blue circles represent the hidden values for benign queries.
Appendix E Safety prompt
We illustrated the safety prompt of LLaMa-2 Official and Self-Reminder in Figure 7.
Figure 7: Illustration of safety prompt used in LLaMa-2 Official and Self-Reminder.
Appendix F LLMs can distinguish harmful queries and harmless queries
Figure 9 visualizes the classifier’s performance on custom and out-of-domain datasets. Table 5 shows the quantitative results of the classifier on different benchmarks. Notably, LLMs demonstrate robust classification capabilities on all datasets except Xstest. On Xstest, a few of benign inputs are interspersed to the harmful side. We speculate that this is because queries in Xstest contains words that are subconsciously harmful but semantically harmless. This observation corresponds to the results that the original LLMs is more prone to rejection on Xstest on Table 1. This poor classification on Xstest aligns with the serious rejections of RDS on Xstest.
Table 5: Evaluation results (AUC) of the classifier on harmful and benign benchmarks. Custom is the training data. ’Others’ includes MaliciousInstruct, AdvBench, and Held-out datasets.
Models Custom Others Xstest
Vicuna-7B-v1.3 1.00 0.99 0.71
Vicuna-13B-v1.3 1.00 0.99 0.83
llama-2-chat-7B 1.00 1.00 0.65
LLaMA3-Instruct-8B 1.00 1.00 0.82
Qwen2-7B-Instruct 1.00 1.00 0.89
(3) i=4
(3) i=5
(3) i=6
(3) i=7
Figure 8: Performance of the classifier at the decoding from the 4-th to 7-th token.
(1) Custom
(2) Out-of-domain benchmarks
(3) Xstest
Figure 9: Performance of the classifier at all datasets. (1) Custom is the training data of the classifier. (2) AdvBench and MaliciousInstruct are the harmful benchmark. Held-out is a benign benchmark. (3) For better visualization, we select symmetrical data from Xstest and visualize both the harmful and benign queries in symmetry pairs.